OpenStudy (anonymous):

A card is selected from a standard deck of 52 playing cards. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). Find the probability of selecting •an odd prime number under 10 given the card is a club. (1 is not prime.) •a Jack, given that the card is not a heart. •a King given the card is not a face card. Show step by step work. Give all solutions exactly in reduced fraction form.

OpenStudy (anonymous):

an odd prime number under 10 given the card is a club 3,5,7 so,3/52 a Jack, given that the card is not a heart. there are 4 jacks in a deck so,3/52 a King given the card is not a face card. King is a face card so,0/52=0

OpenStudy (anonymous):

for the second one wouldn't it be the # jacks (not hearts) be 3 and the total # of not hearts 39 so then it would be 3 / 39 = 1/3 ? please help me

OpenStudy (anonymous):

and the probability (kings) = # kings given which is zero divided by total # not face cards is 52 - 12 wich is 40 then i would have 0 / 40 = no chance

OpenStudy (anonymous):

i am totally lost on this one

myininaya (myininaya):

this was my thinking

myininaya (myininaya):

a) 3/13 b) 1/13 c) 0